For this question I'm suppose to complete the sqare and find the roots of the quadratic equation. Any help would be GREAT!!!
x2 - 2/3x = 0

Answer by stanbon(75887)

complete the sqare and find the roots of the quadratic equation.
x^2 - 2/3x = 0
----------------
Roots:
x(x-2/3)=0
x=0 or x=2/3
----------------------
Complete the square:
x^2-(2/3)x+(1/3)^2)=(1/3)^2
(x+(1/3))^2=(1/3)^2
What do you want to do with this?
Cheers,
Stan H.